OpenStudy (anonymous):

Solve the triangle. A = 32°, a = 19, b = 12

hartnn (hartnn):

do you know sine law ?

hartnn (hartnn):

\(\huge \dfrac{\sin B}{b}=\dfrac{\sin A}{a}\) you can find B from here.

hartnn (hartnn):

yes, can you find B ?

hartnn (hartnn):

\(\huge {\sin B}=\dfrac{\sin A}{a} \times b\) just plug in values... you'll get sin B, and then take arcsin of that to get B.

Parth (parthkohli):

Too small!